# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ 4 Optimal Equilibrium
An equilibrium $\hat{\pi}\in\Pi^{e}_{d}$ is called optimal if $J(\hat{\pi};0,w)\geq J(\pi;0,w)$ for all $\pi\in\Pi^{e}_{d}$ and for all $w>0$.
Recalling the relationship $J(\pi;t,w)=wJ_{0}(t,\pi)$ in Remark 2.3, finding an optimal equilibrium in $\Pi^{e}_{d}$ is equivalent to finding a $\hat{\pi}\in\Pi^{e}_{d}$ that maximizes $J_{0}(t,\pi)$.
@@
$J_{0}(t,\pi^{(T_{0})})=\exp\left\{\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y\right\}l\left(\frac{v^{(T_{0})}(t)}{2}\right).$
@@
According to (4.1), in order to study the maximization of $J_{0}(t,\pi^{(T_{0})})$ over $T_{0}\in\mathcal{T}$, we need to analyze the monotonicity of $\mathcal{L}$ and the monotonicity of $v^{(T_{0})}$ with respect to $T_{0}$. These properties are characterized in Lemma 4.2 and Lemma 4.3, respectively.
@@
$\frac{d\mathcal{L}}{dz}=\frac{1}{h(z)}-\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}z}\right]}{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}z}\right]}=\frac{1}{2h(z)}>0\quad\text{for all }z>0.$
@@
Let $T_{0,1},T_{0,2}\in\mathcal{T}$ with $T_{0,1}<T_{0,2}$. For $t\in[T_{0,2},T]$, we have $v^{(T_{0,1})}(t)=v^{(T_{0,2})}(t)=0$ based on Definition of $\mathcal{T}$. For $t\in[T_{0,1},T_{0,2})$, we have $v^{(T_{0,2})}(t)>0=v^{(T_{0,1})}(t)$. For $t\in[0,T_{0,1})$, we have
@@
By the definition of $\mathcal{T}$, $\Lambda(T_{0,1})>\Lambda(T_{0,2})$. Because $H^{-1}$ is strictly increasing, it follows that $v^{(T_{0,1})}(t)<v^{(T_{0,2})}(t)$ for $t\in[0,T_{0,1})$. Thus, (4.2) is proved.
@@
Under the conditions of Assertion (2) of Theorem 3.2, we have the followings:
@@
Combining (4.1) with Lemma 4.2, we see that $J_{0}(0,\pi^{(T_{0})})$ attains its maximum precisely when $v^{(T_{0})}(0)$ is maximal. Moreover, by Lemma 4.3, $v^{(T_{0})}(0)$ is strictly increasing in $T_{0}\in\mathcal{T}$. Thus, finding the optimal equilibrium strategy is equivalent to identifying the maximal parameter $T_{0}\in\mathcal{T}$. By Lemma 3.3, $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$. Because $\varphi$ is strictly decreasing on $\mathcal{A}$, the existence of a maximal element in $\mathcal{T}$ is equivalent to the existence of a minimal element in $\mathcal{A}$.
If $H(\infty)>\Lambda(0)$, then $\mathcal{A}=[0,\Lambda(0)]$ includes its minimum at 0, and hence a maximal strategy exists. Conversely, if $H(\infty)\leq\Lambda(0)$, then $\mathcal{A}=(\Lambda(0)-H(\infty),\Lambda(0)]$ is left-open and admits no minimum. Thus, $\mathcal{T}$ has a maximal element if and only if $H(\infty)>\Lambda(0)$. In conclusion, the optimal strategy exists if and only if $H(\infty)>\Lambda(0)$, and when it exists, it is given by $\pi^{(\varphi(0))}$.
We now turn to the uniformly optimal equilibrium. In view of the conclusion in assertion (1), we restrict attention to the case $H(\infty)>\Lambda(0)$. Once this condition holds, and following the same line of reasoning as in assertion (1), we immediately obtain that $\pi^{(\varphi(0))}$ is also uniformly optimal, because (4.1) remains valid for every $t\in[0,T]$.
Finally, we consider the uniformly strictly optimal equilibrium. If $\Lambda(t)=0$ for some $t\in[0,T)$, then $v^{(T_{0})}(t)\equiv 0$ for any $T_{0}\in\mathcal{T}$, which implies $J(t,\pi^{(T_{0})})\equiv\exp(\mathcal{L}(0))=1$ for all $\pi^{(T_{0})}\in\Pi_{d}^{e}$. Hence, no uniformly strictly optimal equilibrium exists in this case. Conversely, assume $\Lambda(t)>0$ for all $t\in[0,T)$ and $H(\infty)>\Lambda(0)$. Because $\Lambda(t)>0$ for all $t<T$, we have $\varphi(0)=T$. Let $T_{0}\in\mathcal{T}\setminus\{T\}$, by Lemma 4.3, for any $t\in[0,T)$, it holds that
@@
Because $\mathcal{L}(z)$ is strictly increasing in $z$, we obtain $J_{0}(t,\pi^{(T)})>J_{0}(t,\pi^{(T_{0})})$. Therefore, $\pi^{(\varphi(0))}=\pi^{(T)}$ is the unique uniformly strictly optimal equilibrium. ∎ Analogous to Proposition 3.8, we now investigate the impact of future market opportunities on the risk exposure vector in the case of $\tilde{\mathbb{E}}[\bm{R}]=\infty$. Given the multiplicity of equilibria in this setting, we focus specifically on the optimal equilibrium characterized in Theorem 4.4. The following proposition presents the corresponding comparative statics results. We omit the proof as it follows a similar line of reasoning to that of Proposition 3.8.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}},$
    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
    $\lim_{n\to\infty}a_{n}(0)=\mathbf{0}.$
    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
    $|a_{n}(0)|\searrow 0.$
Here, we adopt the convention that $\frac{1}{0}=\infty$.
@@ 5 Comparative Statics
In the previous sections, we fully characterized the equilibrium strategies associated with a given distribution of risk aversion $\bm{R}$ and discussed how to select the optimal equilibrium. We now turn to comparative statics and ask a natural economic question: how does a shift in the distribution of risk aversion affect the resulting equilibrium investment behavior? In this section, we only consider the case where the equilibrium exists.
To formalize such comparisons, we begin by recalling the notion of first-order stochastic dominance, which provides a standard criterion for ranking risk aversion distributions.
@@
Let $\bm{R_{1}}$ and $\bm{R_{2}}$ be two random variables. We say that $\bm{R_{1}}$ dominates $\bm{R_{2}}$ in the sense of first-order stochastic dominance, denoted as $\bm{R_{1}}\succeq_{1}\bm{R_{2}}$, if
$\tilde{\mathbb{P}}(\bm{R_{1}}\geq x)\geq\tilde{\mathbb{P}}(\bm{R_{2}}\geq x)\quad\text{for all }x\in\mathbb{R}.$
Intuitively, one might expect that if Investor 1 is “more risk-averse” than Investor 2 in the sense of first-order stochastic dominance, then Investor 1 should take a smaller position in the risky asset at all times. Surprisingly, this intuition turns out to be false, as the following numerical example shows.
@@
Assume that there is one stock whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.4$. The time horizon $T=20$. The risk-aversion distributions of the two investors are:
Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$.
Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
@@
Obviously, $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$. Figure 1 displays the comparative evolution of $|a_{i}|$, i=1,2i=1,2.555Because $\bar{\pi}=(\sigma^{\top})^{-1}a$, in the one-dimensional case comparing $|a_{i}(t)|$ is equivalent to comparing $|\pi|_{i}$. More generally, $|a|$ represents the volatility of the investor’s wealth process and thus captures the level of risk borne by the investor. Therefore, in what follows, we focus conceptually on comparing the risk exposure magnitude. The orange-shaded Reversal Region highlights that the more risk-averse Investor 1 adopts a strictly more aggressive position than Investor 2 when the time to maturity is sufficiently long. This counterexample shows that under first-order stochastic dominance, the magnitudes of equilibrium investment strategies need not be monotone.
@@
Figure 1: Comparative evolution of $a_{i}(t)$ under first-order stochastic dominance. The red solid line corresponds to Investor 1 (more risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$; the blue dashed line corresponds to Investor 2 (less risk-averse), with distribution $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$, $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$. The orange shaded region indicates where $|a_{1}(t)|>|a_{2}(t)|$. Parameters: $\lambda=0.4$, horizon $T=20$.
@@
The appearance of this counterexample (Figure 1) may suggest that the search for a general comparative result under first-order stochastic dominance is fundamentally obstructed. In fact, the observed reversal is not an artifact of a particular parameter choice. We show below that such a reversal in investment magnitudes is unavoidable when the investment horizon $T$ is sufficiently large, provided that the risk-aversion distributions satisfy the conditions stated in the following proposition.
In the following proposition, for convenience, we extend the market to an infinite horizon and assume that the market price of risk $\lambda$ is defined on $0,\infty)$. For every fixed $T>0$, we require $\Lambda(0,T)\triangleq\int_{0}^{T}|\lambda(s)|^{2}\mathrm{d}s<\infty$. In addition, we assume that $\lim_{T\to\infty}\Lambda(0,T)=\infty$ and $\lambda(0)\neq 0$. As discussed in the preceding sections, for each fixed $T$, we can define the equilibrium strategy $\{\pi_{i}(t,T):t\in[0,T)\}$ associated with the distribution $\bm{R}_{i}$ (as used in the proposition below). The proposition shows that, as long as the distributions $\bm{R}_{i}$ satisfy certain conditions, the magnitudes $|\pi_{i}(0,T)|$ must exhibit a strict and deterministic ordering when the horizon $T$ is sufficiently large.
@@
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with finite expectations i.e., $\tilde{\mathbb{E}}[\bm{R}_{i}]<\infty$, i=1, 2i=1,\ 2. Assume that the following conditions are satisfied:
  1. (1)
    Let $r_{0}=\operatorname{essinf}\bm{R}_{1}=\operatorname{essinf}\bm{R}_{2}>0$, and assume that there is a positive probability mass at $r_{0}$, i.e., $p_{i}=\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})\in(0,1)$ for $i=1,2$.
  2. (2)
    Define the infimum of the first support point of $\bm{R}_{i}$ strictly greater than $r_{0}$ as $\tilde{r}_{i}=\inf\{r\in supp(\bm{R}_{i}):r>r_{0}\}$, and let $\delta_{i}=\tilde{r}_{i}-r_{0}$ with $\delta_{1}>\delta_{2}>0$.
Under these conditions, for a sufficiently large $T$, we have
$|\pi_{1}(0,T)|>|\pi_{2}(0,T)|.$
@@
As investigated by Desmettre2023, we now focus on the two-point distribution, which represents the closest deviation from the standard expected utility model. This setting allows us to derive a single crossing property, rigorously characterizing the transition from short-term dominance to the long-term reversal observed in Proposition 5.3.
@@
Under the conditions of Proposition 5.3, suppose further that $\lambda(t)\neq 0$ for all $t\in[0,T)$, and that $\bm{R}_{1}$ and $\bm{R}_{2}$ follow two-point distributions with $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$. Moreover, assume that $T$ is sufficiently large so that $|a_{1}(0)|>|a_{2}(0)|$. Then $|a_{1}(\cdot)|$ and $|a_{2}(\cdot)|$ exhibit a single-crossing property: there exists a unique $t^{*}\in[0,T]$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
@@
Having established the single crossing property, we investigate how the probability distribution drives the timing of this reversal. Intuitively, a larger probability assignment accelerates the strategy’s convergence to the log-utility benchmark, leading to a faster intersection as time recedes from the horizon.
@@
Under the conditions of Proposition 5.3 and Proposition 5.4, assume additionally that $p_{1}=p_{2}=p$. Let $t^{*}$ be the intersection time where $|a_{1}(t^{*})|=|a_{2}(t^{*})|$. Then, $t^{*}$ is strictly increasing with respect to $p$.

### v3

@@ 4 Optimal Equilibrium
An equilibrium $\widehat{\pi}\in\Pi^{e}_{d}$ is called optimal if $J(\widehat{\pi};0,w)\geq J(\pi;0,w)$ for all $\pi\in\Pi^{e}_{d}$ and for all $w>0$.
Recalling the relationship $J(\pi;t,w)=wJ_{0}(t,\pi)$ in Remark 2.3, finding an optimal equilibrium in $\Pi^{e}_{d}$ is equivalent to finding a $\widehat{\pi}\in\Pi^{e}_{d}$ that maximizes $J_{0}(t,\pi)$.
@@
$J_{0}(t,\pi^{(T_{0})})=\exp\left(\int_{0}^{v^{(T_{0})}(t)}\frac{1}{h(y)}\mathrm{d}y\right)l\left(\frac{v^{(T_{0})}(t)}{2}\right).$
@@
According to (4.1), in order to study the maximization of $J_{0}(t,\pi^{(T_{0})})$ over $T_{0}\in\mathcal{T}$, we need to analyze the monotonicity of $\mathcal{L}$ and the monotonicity of $v^{(T_{0})}$ w.r.t. $T_{0}$. These properties are provided in Lemma 4.2 and Lemma 4.3, respectively.
@@
$\frac{d\mathcal{L}}{dz}(z)=\frac{1}{h(z)}-\frac{1}{2}\frac{\tilde{\mathbb{E}}\left[\bm{R}e^{-\frac{1}{2}\bm{R}z}\right]}{\tilde{\mathbb{E}}\left[e^{-\frac{1}{2}\bm{R}z}\right]}=\frac{1}{2h(z)}>0\quad\text{for all }z>0.$
@@
Let $T_{0,1},T_{0,2}\in\mathcal{T}$ with $T_{0,1}<T_{0,2}$. For $t\in[T_{0,2},T]$, we have $v^{(T_{0,1})}(t)=v^{(T_{0,2})}(t)=0$ based on the definition of $\mathcal{T}$. For $t\in[T_{0,1},T_{0,2})$, we have $v^{(T_{0,2})}(t)>0=v^{(T_{0,1})}(t)$. For $t\in[0,T_{0,1})$, we have
@@
By the definition of $\mathcal{T}$, $\Lambda(T_{0,1})>\Lambda(T_{0,2})$. Because $H^{-1}$ is strictly increasing, we have $v^{(T_{0,1})}(t)<v^{(T_{0,2})}(t)$ for $t\in[0,T_{0,1})$. Thus, (4.2) is proved.
@@
Under the conditions of Assertion (2) of Theorem 3.2, we have the following assertions.
@@
Combining (4.1) with Lemma 4.2, we see that $J_{0}(0,\pi^{(T_{0})})$ attains its maximum precisely when $v^{(T_{0})}(0)$ is maximal. Moreover, by Lemma 4.3, $v^{(T_{0})}(0)$ is strictly increasing in $T_{0}\in\mathcal{T}$. Thus, finding the optimal equilibrium strategy is equivalent to identifying the maximal parameter $T_{0}\in\mathcal{T}$. By Lemma 3.3, $\mathcal{T}=\{\varphi(\eta):\,\eta\in\mathcal{A}\}$. The existence of a maximal element in $\mathcal{T}$ is equivalent to the existence of a minimal element in $\mathcal{A}$, because $\varphi$ is strictly decreasing on $\mathcal{A}$.
If $H(\infty)>\Lambda(0)$, then $\mathcal{A}=[0,\Lambda(0)]$ includes its minimum at 0, and hence a optimal strategy exists. Conversely, if $H(\infty)\leq\Lambda(0)$, then $\mathcal{A}=(\Lambda(0)-H(\infty),\Lambda(0)]$ is left-open and admits no minimum. Thus, $\mathcal{T}$ has a maximal element if and only if $H(\infty)>\Lambda(0)$. In conclusion, the optimal strategy exists if and only if $H(\infty)>\Lambda(0)$, and when it exists, it is given by $\pi^{(\varphi(0))}$.
Following the same line of reasoning as in Assertion (1), we immediately obtain that $\pi^{(\varphi(0))}$ is also uniformly optimal as (4.1) remains valid for every $t\in[0,T]$.
Assume that $\Lambda(t)=0$ for some $t\in[0,T)$. Then $v^{(T_{0})}(t)\equiv 0$ for any $T_{0}\in\mathcal{T}$, which implies $J(t,\pi^{(T_{0})})\equiv\exp(\mathcal{L}(0))=1$ for all $\pi^{(T_{0})}\in\Pi_{d}^{e}$. Hence, there is no uniformly strictly optimal equilibrium in this case. Conversely, assume that $\Lambda(t)>0$ for all $t\in[0,T)$ and $H(\infty)>\Lambda(0)$. We have $\varphi(0)=T$ because $\Lambda(t)>0$ for all $t<T$. Let $T_{0}\in\mathcal{T}\setminus\{T\}$. By Lemma 4.3, for any $t\in[0,T)$, it holds that
@@
We obtain $J_{0}(t,\pi^{(T)})>J_{0}(t,\pi^{(T_{0})})$ as $\mathcal{L}(z)$ is strictly increasing in $z$. Therefore, $\pi^{(\varphi(0))}=\pi^{(T)}$ is uniformly strictly optimal. ∎
In analogy with Proposition 3.8, we now analyze the effect of future market opportunities on the risk exposure vector for the case $\tilde{\mathbb{E}}[\bm{R}]=\infty$. Given the multiplicity of equilibria in this setting, we focus on the optimal equilibrium characterized in Theorem 4.4. The following proposition presents the corresponding results. We omit the proof as it follows a similar line of reasoning to that of Proposition 3.8.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}$, where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}$ as $n\to\infty$.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then $\lim\limits_{n\to\infty}a_{n}(0)=\mathbf{0}$. Moreover, $|a_{n}(0)|\searrow 0$ as $n\to\infty$.
Here we adopt the convention $\frac{1}{0}=\infty$.
@@ 5 Comparative Statics
In the previous sections, we have derived, in closed form, the equilibrium strategies associated with a given distribution of the RRA $\bm{R}$ and discussed how to select the optimal equilibrium. We now turn to comparative statics and consider a natural question: how does a shift in the distribution of the RRA affect the resulting equilibrium investment behavior?
Throughout this section, we write $H_{i}$ for $H_{{\bm{R}_{i}}}$ and $a_{i}$ for $a_{{\bm{R}_{i}}}$, $i=1,2$.
5.1 Two-Point Distributed RRA and Single-Crossing
We begin by considering the first-order stochastic dominance, which should be a natural way to rank the risk aversion distributions.
@@ 5.1 Two-Point Distributed RRA and Single-Crossing
Let $\bm{R}_{1}$} and $\bm{R}_{2}$} be two random variables. We say that $\bm{R}_{1}$} dominates $\bm{R}_{2}$} in the sense of first-order stochastic dominance, denoted ${\bm{R}_{1}}\succeq_{1}{\bm{R}_{2}}$, if $\tilde{\mathbb{P}}({\bm{R}_{1}}\geq x)\geq\tilde{\mathbb{P}}({\bm{R}_{2}}\geq x)$ for all $x\in\mathbb{R}$.
Intuitively, one might expect that if ${\bm{R}_{1}}\succeq_{1}{\bm{R}_{2}}$, then the corresponding risk exposure magnitudes $|a_{i}(\cdot)|$ ($i=1,2$) should satisfy $|a_{1}(t)|\leq|a_{2}(t)|$ for any $t\in[0,T)$. Surprisingly, this intuition turns out to be false, as the next example with two-point distributed RRA shows.
@@
Assume that there is only one stock, whose market price of risk is constant, $\lambda(\cdot)\equiv\lambda=0.4$. The time horizon $T=20$. There are two investors with distributions of RRA:
  • Investor 1: $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$;
  • Investor 2: $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$.
@@
Obviously, $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$. Figure 1 displays the comparative evolution of $|a_{i}(\cdot)|$ (i=1,2i=1,2). [^4] Because $\bar{\pi}=(\sigma^{\top})^{-1}a$, in the one-dimensional case comparing $|a_{i}(\cdot)|$ is equivalent to comparing $|\pi_{i}(\cdot)|$. More generally, $|a|$ represents the local volatility of the investor’s wealth process and thus captures the level of risk borne by the investor. Therefore, in what follows, we focus conceptually on comparing the risk exposure magnitude (local volatility) instead of the trading strategy. The red-shaded Reversal Region highlights that Investor 1 adopts a strictly riskier position than Investor 2 when the time to maturity is sufficiently long. This counterexample shows that a “larger” RRA in the sense of first-order stochastic dominance does not necessarily yield a less risky investment.
@@
Figure 1: Comparative evolution of $|a_{i}(\cdot)|$ ($i=1,2$) under the first-order stochastic dominance. The red solid line corresponds to Investor 1 with $\tilde{\mathbb{P}}(\bm{R}_{1}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{1}=3)=0.1$; the blue dashed line corresponds to Investor 2 with $\tilde{\mathbb{P}}(\bm{R}_{2}=1)=0.9$ and $\tilde{\mathbb{P}}(\bm{R}_{2}=2)=0.1$. The red shaded region indicates where $|a_{1}(\cdot)|>|a_{2}(\cdot)|$. Parameters: $\lambda=0.4$, $T=20$.
@@
The appearance of Example 5.2 might suggest that establishing a general comparative result under the first-order stochastic dominance is fundamentally obstructed. In fact, the observed reversal is not merely an artifact of a particular parameter choice.
In the following proposition, we focus on the case where the risk aversions $\bm{R}_{i}$ ($i=1,2$) satisfy $\operatorname{essinf}\bm{R}_{1}=\operatorname{essinf}\bm{R}_{2}=r_{0}>0$. In this case, $H_{i}(\infty)=\infty$ ($i=1,2$) (see Lemma A.2). Consequently, according to Theorem 3.2(1), for any deterministic and right-continuous market price of risk $\lambda(\cdot)$ in $L^{2}([0,T))$, the equilibrium strategy exists and is unique.
The next proposition shows that, under suitable conditions on $\bm{R}_{i}$, in particular when $\bm{R}_{1}$ and $\bm{R}_{2}$ share the same “smallest” value but $\bm{R}_{1}$ has a larger “second-smallest” value than $\bm{R}_{2}$, there exists $\overline{\Lambda}>0$ such that, for any $\lambda(\cdot)$ with $\lambda(0)\neq 0$ and $\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$, the corresponding risk exposure magnitudes satisfy $|a_{1}(0)|>|a_{2}(0)|$. Moreover, if $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda(T-)\neq 0$, then $|a_{1}(T-)|<|a_{2}(T-)|$.
@@
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables with $\tilde{\mathbb{E}}[\bm{R}_{i}]<\infty$ $(i=1,2)$ satisfying the following two conditions.
  1. (c1)
    Let $r_{i}\triangleq\operatorname{essinf}\bm{R}_{i}$, $i=1,2$. Assume that $r_{0}\triangleq r_{1}=r_{2}>0$ and $p_{i}\triangleq\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})\in(0,1)$ for $i=1,2$.
  2. (c2)
    Let $\tilde{r}_{i}\triangleq\inf\{r\in\operatorname*{supp}(\bm{R}_{i}):r>r_{0}\}$ and $\delta_{i}\triangleq\tilde{r}_{i}-r_{0}$, $i=1,2$. Assume that $\delta_{1}>\delta_{2}>0$.
Then, there exists $\overline{\Lambda}>0$ such that, for any $\lambda(\cdot)$ with $\lambda(0)\neq 0$ and $\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$, the corresponding risk exposure magnitudes satisfy $|a_{1}(0)|>|a_{2}(0)|$. In addition, if $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda(T-)\neq 0$, then $|a_{1}(T-)|<|a_{2}(T-)|$.
@@
Next, under the conditions of Proposition 5.3, we further investigate the crossing properties of the risk exposure magnitudes $|a_{i}(\cdot)|$ ($i=1,2$) corresponding to the two distributions. This includes the single-crossing property (Proposition 5.4) as well as the monotonicity of the crossing time w.r.t. the distribution parameters (Propositions 5.6 and 5.7). To this end, we focus on the two-point distributed RRA, which, as investigated by Desmettre and Steffensen (2023), is the closest RRA model to the standard expected utility framework. Specifically, we assume that the risk aversion $\bm{R}_{i}$ ($i=1,2$) satisfies
$\begin{split}&\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0})=p_{i}=1-\tilde{\mathbb{P}}(\bm{R}_{i}=r_{0}+\delta_{i})\quad\text{with }r_{0}>0,\,p_{i}\in(0,1),\delta_{1}>\delta_{2}>0.\end{split}$ (5.1)
Below, we begin by establishing the single-crossing property of the risk exposure magnitudes; that is, under certain conditions, there exists a unique time $t^{*}$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*})|$.
@@
Suppose that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) and $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$. [^5] The condition $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ is equivalent to $(1-p_{1})\delta_{1}>(1-p_{2})\delta_{2}$. Let $\widehat{\Lambda}\in(0,\infty)$ be the unique solution to the equation
[…]
Then, for any $\lambda(\cdot)$ satisfying $\lambda(t)\neq 0$ for all $t\in[0,T)$ and $\lambda(T-)\neq 0$, we have the following assertions:
  1. (1)
    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\widehat{\Lambda}$, then $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in[0,T)$.
  2. (2)
    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t>\widehat{\Lambda}$, then there exists a unique $t^{*}\in(0,T)$ such that $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in(t^{*},T)$ and $|a_{1}(t)|>|a_{2}(t)|$ for all $t\in[0,t^{*})$.
  3. (3)
    If $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t=\widehat{\Lambda}$, then $|a_{1}(0)|=|a_{2}(0)|$ and $|a_{1}(t)|<|a_{2}(t)|$ for all $t\in(0,T)$.
@@
Remark 5.5.
It is important to note that the conditions in Propositions 5.3 and 5.4 are compatible with first-order stochastic dominance. One can construct risk aversion distributions such that $\bm{R}_{1}\succeq_{1}\bm{R}_{2}$ holds while the conditions in Propositions 5.3 and 5.4 are simultaneously satisfied: letting $p_{1}=p_{2}$ in (5.1). For instance, the parameters in Example 5.2 satisfy these conditions with $r_{0}=1$, $\delta_{1}=2$, $\delta_{2}=1$, $\tilde{\mathbb{E}}[\bm{R}_{1}]=1.2>1.1=\tilde{\mathbb{E}}[\bm{R}_{2}]$ and $\lambda\equiv 0.4\neq 0$. In particular, Figure 1 illustrates both the reversal in risk exposure magnitude and the single-crossing property.
Having established the single-crossing property, we next investigate the monotonicity of $t^{*}$ w.r.t. $p\triangleq p_{1}=p_{2}$ under the conditions of Proposition 5.4.
@@
Let $T>0$ be fixed. Assume that the risk aversions $\bm{R}_{1}$ and $\bm{R}_{2}$ satisfy (5.1) with $p_{1}=p_{2}=p\in(0,1)$ and that $\lambda(t)\neq 0$ for all $t\in[0,T)$ with $\lambda(T-)\neq 0$. Let $a_{i,p}$ denote the risk exposure vector corresponding to $\bm{R}_{i}$ with $p_{i}=p$ for $i=1,2$. Let $t^{*}(p)\in(0,T)$ be the crossing time, where $|a_{1,p}(t^{*}(p))|=|a_{2,p}(t^{*}(p))|$. Then the set
$I\triangleq\left\{p\in(0,1):t^{*}(p)\in(0,T)\right\}$
is open and the crossing time $t^{*}(\cdot)$ is continuously differentiable in II. [^6] The set $I$ is non-empty if $\Lambda(0)$ is sufficiently large. Indeed, for any fixed $p\in(0,1)$, Proposition 5.4 ensures that the crossing time $t^{*}(p)$ exists in $(0,T)$ provided that $\Lambda(0)$ exceeds the threshold $\widehat{\Lambda}$ defined therein, which depends on $p$. Moreover, its derivative satisfies
$\frac{\mathrm{d}t^{*}}{\mathrm{d}p}(p)>0\quad\text{for all }p\in I.$